Subject: 3.5-m Users Committee minutes 11/27/06
From: Bruce Gillespie
Submitted: Sat, 2 Dec 2006 16:50:47 -0700
Message number: 1000
(previous: 999,
next: 1001
up: Index)
APO 3.5-m Users Committee Phonecon, 11/27/06
Attending: Suzanne Hawley, Bruce Gillespie, Michael Strauss, Scott
Anderson, Jon Fulbright, John Bally, Al Harper, Remy Indebetouw, Rene
Walterbos
Absent: Jon Holtzman, Russet McMillan, Don York
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael said the PU users are concerned about
DIS--the red camera on DIS is currently out of service. This was
discussed at some length later in the meeting.
JHU (Fulbright) - Jon mentioned that one of the JHU users wondered
why we don't use Skype between the observatory and the remote users.
After some group discussion, we concluded that what we have now (text
windows, regular phones) is adequate communication for routine
operations, but that a video conferencing system could be useful for
troubleshooting problems.
Chicago (Harper) - Al said there was nothing new, but he was
interested in hearing about the NIC-FPS Fabry-Perot situation.
Washington (Anderson) - Scott thanked Suzanne for getting the DIS and
NIC-FPS status reports to the affected users so promptly.
NMSU (Walterbos) - Rene mentioned that he recently used DIS in blue-
only mode, and that the blue spectrum looks wonderful--no stray
light, good signal strength.
Virginia (Indebetouw) - Remy had no comments.
Colorado (Bally) - John reported that tests of the recently observed
gradient in NIC-FPS show that it can be removed during data
reduction, and that the instrument's performance is fairly nominal.
The Fabry-Perot etalon is still in England, where the controller is
being modified to give it a longer time constant. The present hope is
that the Fabry-Perot might be re-installed during next summer's
shutdown.
**********************************
Discussion of telescope/instruments report:
The report is given below; there was considerable discussion about
the DIS upgrades and the situation with the red CCD. Since the
report was written, a plan has been finalized to bring Jim Gunn, Dale
Sandford, and Connie Rockosi to APO December 8 - 11 to retry getting
the new chip working, or else to reinstall the old red chip. In the
meantime, the DIS red camera is unavailable but the DIS blue camera
is working well.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
10/14/06 through 11/22/06
Bruce Gillespie
0) Overview
After an earlier-than-usual snowstorm in mid-October, we've enjoyed
unseasonably warm and clear weather throughout the period. The
telescope performance was very good, and we are seeing additional
indications of pointing and imaging improvements, probably
attributable to the new top end. A change in the performance of NIC-
FPS was noticed and characterized, and "DIS-fest" occurred with mixed
success.
1) Telescope
Aside from enjoying the largely excellent on-sky telescope
performance, we did some engineering work on the new 2ndary
transverse actuators, and additional characterization tests of the
NA2 rotator. With input from John Wilson (UVa), Mike Carr
(Princeton), and Larry Carey (UW), a recommendation for a rotator
upgrade path was made to Mark Klaene (APO), one that we decided to
implement in 2007 as a relatively small CIF project. Also, a test
version of an upgraded TUI (V. 1.3a9) was released for site testing;
this version has improvements in the focus scripts for DIS and the
echelle.
2) Instruments
After the end of the summer shutdown, it was noticed the NIC-FPS raw
images had a new and unexpected density gradient near one edge of the
chip. After discussions with Fred Hearty and Stephane Beland (CU)
who were largely responsible for building and commissioning the
instrument, we learned that this is a reprise of a much more serious
problem that was discovered and corrected during commissioning. CU
personnel obtained some observations and verified that the gradient
should be removable during data reduction. Stephane is planning to
re-tweak the electronics in January, and in the meantime we have
alerted the NIC-FPS users to be aware of the situation and to be
extra careful about their data reductions.
As planned, we took DIS out of service for a week in early November
to complete elements of the DIS upgrade project. Jim Gunn and Mike
Carr (Princeton), Dale Sandford (Yerkes), and Jeff Morgan (UHawaii)
were assisted by site staff in what was planned to be the
installation of the new blue prism assembly, replacement of the field
lens in the red dewar, and replacement of the red CCD with a deep-
depletion device. Everything was accomplished, not without
considerable troubles, with the exception of replacing the red CCD.
Among the unexpected troubles were CryoTiger-caused contamination in
both the blue and red dewars (which were then cleaned), trouble with
the slit wheel mechanism during reinstallation of the blue prism, and
the inability of getting the new CCD to act like a CCD, largely due
to confusion and misinformation from the chip's vendor on how it
operates (we had been led to believe it would be "plug and play").
Our "dream team" of experts ran out of time trying to get the chip to
work correctly, and left the site flummoxed and exhausted. Before
taking additional steps, we decided to regroup. We now plan to
reinstall the old chip at the earliest opportunity (tentatively
during the second week in December). In the meantime, users have
been told to use DIS blue-only, or to find substitute observing
programs. Reinstallation of the new red chip will be contingent on
obtaining a complete understanding of its operating properties.
The good news on DIS is that there are preliminary indications that
the blue sensitivity of DIS is wonderful and as it was meant to be;
the increase in throughput at 3700 Angstroms was measured as ~3x
better than with the old prism, and is yet even higher at shorter
wavelengths.
The Goddard Fabry-Perot is temporarily out of service while the F-P
instrument team effects some repairs and upgrades. SPIcam is having
trouble keeping its cryo-temperature stable, and will be taken out of
service at a convenient time to replace its failing ion pump.
3) Engineering and CIF projects
In the past month, we have continued planning of the CIF work list
for next year; we worked on cleanup, characterization, and
documentation for the new top end; ran DIS-fest (see above); planned
for the NA2 rotator upgrade for TripleSpec; started thinking about
auto-focus; and continued design work for the telescope motor
controllers upgrade.
4) Miscellaneous
Nick McDonald has accepted our offer to become Telescope Systems
Engineer and will start in January, filling the post recently vacated
by Jon Davis. We have also posted a recruitment notice for an
Electronics Engineer, a position that will be shared with the SDSS-II
project.
**********************************
Future of old top end - Hawley:
Since no one has offered a plan to reuse the old top end, we have
decided to scrap it. Rene asked Bruce if the future of the old top
end had been discussed with Kurt Anderson and Bruce said he'd bring
it up with Kurt on his next site visit.
**********************************
New computer and security procedures - Hawley/Gillespie
Suzanne said that we have a new and powerful data reduction computer
at APO, named Newton (replacing Tycho). Fritz Stauffer has
implemented security protocols for this machine (see below), which
will require quarterly password changes. These passwords, specific
to each institution, will be sent to the Users Committee members at
each institution by snail mail, with an e-mail reminder, for both the
current quarter and Q1/07. John Bally asked if sftp is an allowable
transfer protocol. Bruce checked on this later with Fritz, and found
that sftp is allowed. Note that scp is also available for file
transfer to/from Newton.
**********************************
BoG presentations, decisions, annual reports - Hawley:
Suzanne reported that the ARC annual Board of Governors meeting was
held in Chicago on 13 November. The presentation materials are
posted at a blind URL that the Users Committee members have been
sent. The board members seemed quite pleased with the breadth and
quantity of science and engineering accomplishments in the past year,
and approved our budget request for 2007. New in next year's budget
is a staff position (to be shared with SDSS) for an electronics
engineer. Rene asked if we had sent the '05/06 publications list to
the Users Committee. Bruce said no, but we can load it on the APO
website.
ACTION: Rene to collect history of 3.5-m publication, and Bruce will
make a space for them on the APO website.
There followed a discussion of the post-2008 plans for the 2.5-m, and
the potential impact to the 3.5-m operations. Bruce agreed to send a
copy of the post-2008 recommendations to the users committee members
who may discuss with interested people at their institutions (done
11/28).
**********************************
Systems workshop - Hawley:
Suzanne reported that she had recently attended the 3rd Observatory
Telescope Systems Workshop, held in Scottsdale, AZ in mid-November.
This is an AURA-led activity that promotes the concept of the public
and private observatories working together in a coordinated manner.
The TSIP program grew out of this, but is generally only effective
for the larger telescopes--too expensive for the small and mid-sized
telescopes to use. The PREST program was devised to help small
telescopes, in the 1- to 2-meter class. During this 3rd meeting, it
was discussed how to get more money for these programs, and there was
some push to find programs that would support the mid-sized
telescopes, especially since the Senior Review has advocated
continuing support of mid-sized telescopes.
Suzanne said her talk on the 3.5-m telescope was well received, and
was striking to the audience in that we are ~90% remote operation,
can do quick instrument changes, respond to ToO opportunities,
support synoptic and time-domain programs, and may soon have NIR
Fabry-Perot capabilities. There followed a discussion about whether
we should write an SPIE paper that describes the 3.5-m telescope
system. Bruce mentioned that he wrote such a paper in 2002 (see
http://www.apo.nmsu.edu/Telescopes/eng.papers/SPIEpapers/
spie2002.pdf), and it is probably time to update it (although he
didn't volunteer to do it anytime soon).
**********************************
Actions from earlier meetings - group
[open from previous months]:
ACTION: Bruce to discuss issues of computer and network security
with Fritz Stauffer. STATUS: Closed, discussed at today's meeting
(see below).
ACTION: Users Committee members should send Michael
(strauss@astro.princeton.edu) the names and e-mail addresses of 3.5-m
users who are not yet signed up for the apo35general mailer.
STAUS: No recent activity reported, closed.
[new actions from last month]:
ACTION: Users Committee representative to gather one or two science
highlights and send them to Bruce and Suzanne not later than mid-
October.
STATUS: Closed.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 22 January, at
8:30 AM Pacific Time. Agenda and other materials will be sent to the
committee members during the preceding week.
-------------------------------------------------------------------
[announcement from Fritz Stauffer on the new data reduction computer
at APO, and new security protocols]:
APO has a new computer, newton.apo.nmsu.edu, to replace tycho. It
has been set up with accounts for each of the ARC members and a
generic account,
visitor1, for all other users. The computer has been set up with
most of the popular astronomy packages.
The accounts have:
10 GB of permanent disk space set by quotas
750 GB of scratch space with files > 1 week automatically
deleted
passwords changed quarterly and mailed to the Users
Committee members
of each institution and kept locally
ssh access
ftp is not allowed due to security risks
ftp to images is still available with the usual (previously known)
password.
tycho will be online for a while to help transition to newton, but
users are encouraged to start using newton.
[note added by SlH: ** See your users committee member to get the
password, and remember
there is a new one at the start of each quarter. **]
Here is a description of the 3.5m users computing facitilites:
http://www.apo.nmsu.edu/35m_operations/35m_manual/
General_Information/PDF/APO-Computing-Handout.pdf
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1000 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO